OpenStudy (anonymous):

3 1/2 times 3/4 is 2.625 correct?

OpenStudy (anonymous):

\[31/2*3/4 = 93/8 = 11.625\] And \[3\frac{ 1 }{ 2 }*\frac{ 3 }{ 4 } = \frac{ 7 }{ 2 }*\frac{ 3 }{ 4 }= \frac{ 21 }{ 8 } = 2.65 \]

OpenStudy (anonymous):

So, yeah you are correct.

OpenStudy (anonymous):

Oh sweet! Thank you!